A magazine survey found that women over the age of 55 consume anaverage of 1660 calories a day. In order to see if the same wastrue for women over 55 in assisted living facilities, a researchersampled 43 women. She found that the mean for the sample was 1446and that the sample standard deviation was 56 calories. Atα=.10 test the claim that the average number of caloriesconsumed by the women is assisted living is the same as the numbercalories consumed by the women magazine survey.
1
Expert's answer
2019-07-29T09:58:42-0400

The provided sample mean is "\overline{X} =1446" and the sample standard deviation is "s=56", and the sample size is "n=43."

The following null and alternative hypotheses need to be tested:


"H_0: \mu=1660""H_1: \mu\not =1660"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.1," and the critical value for a two-tailed test is "t_c=1.682."

The rejection region for this two-tailed test is "R=\{{t: |t|>1.682}\}"  

The t-statistic is computed as follows:


"t={\overline{X}-\mu \over s/\sqrt{n}}={1446-1660 \over 56/\sqrt{43}}=-25.0588"

Since it is observed that "|t|=25.0588>t_c=1.682,"  |t| = 25.059 it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=0," and since "p=0<0.1," it is concluded that the null hypothesis is rejected.

Conclusion

It is concluded that the null hypothesis "H_0" is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is different than 1660, at the 0.1 significance level.


The 90% confidence interval is 

"1431.636<\mu<1460.364"